# Self-Introspective Decoding: Alleviating Hallucinations for Large Vision-Language Models

### v2

@@ Abstract
Hallucination remains a significant challenge in Large Vision-Language Models (LVLMs). To alleviate this issue, some methods, known as contrastive decoding, induce hallucinations by manually disturbing the raw vision or instruction inputs and then mitigate them by contrasting the outputs of the original and disturbed LVLMs. However, these holistic input disturbances sometimes induce potential noise and also double the inference cost. To tackle these issues, we propose a simple yet effective method named Self-Introspective Decoding (SID). Our empirical investigations reveal that pre-trained LVLMs can introspectively assess the importance of vision tokens based on preceding vision and text (both instruction and generated) tokens. Leveraging this insight, we develop the Context and Text-aware Token Selection (CT2S) strategy, which preserves only the least important vision tokens after the early decoder layers, thereby adaptively amplify vision-and-text association hallucinations during auto-regressive decoding. This strategy ensures that multimodal knowledge absorbed in the early decoder layers induces multimodal contextual rather than aimless hallucinations, and significantly reduces computation burdens. Subsequently, the original token logits subtract the amplified fine-grained hallucinations, effectively alleviating hallucinations without compromising the LVLMs’ general ability. Extensive experiments illustrate that SID generates less-hallucination and higher-quality texts across various metrics, without much additional computation cost.111Codes are in the appendix and also available at https://github.com/huofushuo/SID.
@@ 1 Introduction
Despite their extraordinary versatility, LVLMs face a significant challenge known as the ‘hallucination’. Concretely, hallucinated texts are fluent and semantically coherent but contain incorrect or non-exist statements about the given image, e.g., generating irrelevant or meaningless responses, identifying inaccurate colors, numbers, and locations of objects not present in the image (Huang et al., 2024). This flaw in LVLMs poses a significant risk for real-world applications to become trustworthy AI assistants. For instance, in model-assisted computer-aided diagnosis scenarios (Wang et al., 2023), such misinterpretation of medical images could lead to serious medical accidents.
@@ 2 Related Work
Hallucination in Foundation Models. Hallucination, defined as the generation of irrelevant, factually incorrect, or meaningless text in a given context (Rohrbach et al., 2018; Zhang et al., 2024; Guan et al., 2024; Wu et al., 2024b), is a significant bottleneck in current foundation models. This issue can stem from overfitting specific patterns in the training data, a lack of understanding world knowledge, or an inability to effectively contextualize a given input (Ji et al., 2023). In the context of LLMs, hallucinations often manifest as generated content that conflicts with world knowledge or common sense. For LVLMs, the primary concern is whether the generated answer conflicts with the provided images. To mitigate the hallucination issue, several solutions have been proposed, including robust instruction tuning with curated datasets (Lee et al., 2022; Gunjal et al., 2024; Liu et al., 2024a; Zhao et al., 2023; Jiang et al., 2024; Yu et al., 2024b; a; Ma et al., 2024; Yue et al., 2024), post-hoc utilizing auxiliary analysis networks (Manakul et al., 2023; Zhou et al., 2024; Yin et al., 2023; Chen et al., 2024b; Wu et al., 2024a; Feng et al., 2024), and various decoding strategies (Li et al., 2022; Chuang et al., 2024; Liu et al., 2024c; Leng et al., 2024; Favero et al., 2024; Wang et al., 2024a; Kim et al., 2024; Zhu et al., 2024). However, robust instruction tuning requires massive high-quality datasets and advanced GPU clusters, making it resource-intensive; Post-hoc utilizing auxiliary networks heavily rely on the auxiliary network, leading to high inference costs. As for decoding strategies, representative LVLMs hallucination alleviation methods (Leng et al., 2024; Favero et al., 2024; Wang et al., 2024a) manually disturb raw inputs to induce hallucinations and then contrast them to alleviate the issue. However, holistic disturbing raw inputs might bring additional noise during contrastive decoding, and double the inference cost. In this paper, we propose an efficient Self-Introspective Decoding (SID) that induces and then mitigates vision-and-text association hallucination by token-level disturbances, greatly reducing the inference cost.
@@ 3.1 Paradigm of LVLMs Generation
Vision and Language Inputs. The inputs of LVLMs consist of both image () and text (). Generally, the raw images are commonly fed to the visual encoder, and then the cross-model projection module maps vision information into LLMs’ input space, which is denoted as vision tokens ( is the length of vision tokens). Similarily, text is processed by tokenizer and embedding modules, which is denoted as text tokens ( is length of text tokens). Then, the image () and text () tokens are concatenated as the final input of LLMs.
@@ 3.2 Re-thinking Contrastive Decoding in LVLMs
However, we argue that manually disturbing raw inputs might not trigger the desired hallucinations, while holistic disturbances will bring uncertainty noise that compromises the normal decoding. To validate our claim, we analyze the performances of normal decoding, VCD, and ICD using the POPE (Li et al., 2023d) metric, under both sampling and greedy decoding settings. POPE quantitatively converts the hallucination evaluation into a binary classification problem by using the question format to prompt the model: ‘Is there a <object> in the image?’, with expected answers being ‘Yes’ or ‘No’. From Table 1, under the greedy decoding setting, CD methods improve performance in the adversarial setting, which are more challenging as they prioritize co-occurring confusing objects. CD methods achieve this by exacerbating and subtracting hallucinated concepts from the original distribution. However, in random settings, where objects absent from the image are chosen randomly and are easily recognized, CD methods slightly underperform normal greedy decoding, which indicates that the correct token logit is somewhat compromised during contrastive decoding. In the sampling decoding setting, CD methods clearly outperform the normal sampling decoding. However, CD methods rely on the adaptive plausibility constraint (Eq. 3) to filter out low-probability tokens. Without Eq. 3, CD methods are inferior to normal decoding in both random and adversarial settings, validating that vision-and-text agnostic input distributions induces potential uncertainty noise after Eq. 2. To address these issues, we propose a decoding strategy named Self-Introspective Decoding (SID). SID adaptively amplifies vision-and-text association hallucinations informed by generated tokens to guide LVLMs in exploring factualness. Details are illustrated in the Sec. 4 and Fig. 2.
@@ 4 Methodology
Analyses. Fig. 3 and 4 preliminarily valid the efficacy of qualitatively. In Fig. 3, the preserved least important tokens mainly reflect areas opposite to the query. For instance, when querying ‘cup’ in Fig. 3 (left), LVLMs focus on ‘cup’ in the foreground, thus preserving background tokens with low . Conversely, LVLMs pay attention to background items when querying ‘couch’. When querying existing items in Fig. 3 (right), vision tokens of unrelated regions are mainly preserved. For open-end generative tasks in Fig. 4, auto-regressive decoded tokens are generated based on preceding vision (), instruction (), and generated text () tokens. The preserved vision tokens are adaptively adjusted according to preceding tokens at each decoding step, primarily focusing on spurious related regions. More quantitative analyses are provided in Appendix A.6, where Table 8 illustrates that vision tokens with high greatly maintain original ability, while tokens with low reach nearly accuracy in binary classification. Above evaluations suggest that Eq. 5 effectively assesses the importance of vision tokens.
We further demonstrate the open-end generated hallucinations induced by ours, Vision Disturbance (VD) (Leng et al., 2024), and Instruction Disturbance (ID) (Wang et al., 2024a) in Fig. 5 and 10. The hallucinations we amplified are more vision-and-text association compared to VD, while LVLMs usually refuse to response to ID. Additionally, we demonstrate the quantitative results for discrimination and generation tasks with VD and ID as inputs in Appendix Table 8. Interestingly, VD and ID do not degrades much, especially in discrimination tasks. Experiments imply that disturbed target logits still have the highest probability in most cases, and therefore, contrastive decoded target logits are not enhanced much after Eq. 2, while CD methods are susceptible to potential uncertainty noise.
@@
Based on the above investigations, we argue that to induce context- and text-aware hallucinations for contrastive decoding, only a small percentage of vision tokens with low attention scores should be preserved after the early decoder layers. To validate our claims, we conduct the following experiments: 1) In Vision Encoder (VE), we preserve tokens with low attention values between the token and vision tokens in the penultimate layer, calculated as: . 2) In the LLM decoder, we preserve tokens with low importance score (Eq. 5) across varying layers (). Additionally, we adjust the number of preserved vision tokens. As shown in Fig. 6, firstly, pruning vision tokens in VE based on may not always yield positive gains, as the token lacks information about instructions and generated texts, which are crucial for multimodal understanding. Specifically, pruning all vision tokens resembles VIG (Favero et al., 2024), which contrastively amplifies the vision importance over the language prior by ablating vision inputs. Secondly, aggressive pruning vision tokens (i.e., 0) after the is not optimal. The loss of vision information for subsequent decoder layers results in losing the visual contexts, leading to aimless hallucinations without sufficient vision grounding. Thirdly, selecting tokens in the late decoder layers degrades contrastive decoding to normal decoding, as preceding layers of LVLMs already decode and understand multimodal information, which is consistent with LLMs’
@@ 5.1 Experimental Settings
Models and Baselines. We utilize four representative LVLMs: InstructBLIP (Dai & et al, 2023), Shikra (Chen et al., 2023a), LLaVA-1.5 (Liu et al., 2024b) at the 7B scale, and LLaVA-NeXT (Li et al., 2024) at the 8B scale. For detailed model descriptions and results at larger scales, please refer to Appendix A.2 and A.6. Since our method aims to propose training-free LVLM decoding strategies, we compare six decoding methods: Sampling (Top-p=1), Greedy, Dola (Chuang et al., 2024), and LVLM decoding strategies (VCD (Leng et al., 2024), ICD (Wang et al., 2024a), and OPERA (Huang et al., 2024)). For comprehensive comparisons, we apply VCD and ICD in both sampling (Top-p=1) and greedy decoding settings. Additionally, we validate SID with different decoding strategies, as detailed in Appendix A.6.
@@ 5.2 Evaluation Results
In this section, we follow previous methods (Leng et al., 2024; Wang et al., 2024a; Huang et al., 2024) to evaluate the SID on CHAIR (Rohrbach et al., 2018) and POPE (Li et al., 2023d) metrics. Besides manually designed metrics, we also leverage GPT-4 assisted benchmark (Zhao et al., 2023) to evaluate attribute, location, and relation hallucinations. MME (Fu et al., 2023) and MMBench (Liu et al., 2023b) benchmarks are employed to assess the LVLM’s general ability. Moreover, GPT4-V assisted evaluation (Appendix A.4) on both hallucination alleviation and generated text quality and case studies (Appendix A.5) on LLaVA-Bench-in-the-Wild are in Appendix.
@@
CHAIR and POPE Evaluations. CHAIR (Rohrbach et al., 2018) and POPE (Li et al., 2023d) are quantitative metrics to assess objection hallucinations of VLMs. Please refer to the Appendix A.3 for detailed descriptions of CHAIR and POPE. As for CHAIR, Following (Wang et al., 2024a; Huang et al., 2024; Yue et al., 2024), we randomly select 500 images from the validation set of the MSCOCO (Lin et al., 2014) dataset and query different LVLMs with the prompt: ‘Please describe this image in detail.’. We set the max new tokens to 512 to generate responses for fair comparisons. As shown in Table 3, our method outperforms other baselines in most cases, validating the effectiveness of SID in open-end generation tasks. Compared to CD methods, SID online adaptively prunes attention-important vision tokens informed by instruction and generated text to induce fine-grained hallucinations for contrastive decoding during open-end text generations. For the POPE metric, which comprises three datasets, we average the results in Table 4. Our method performs best overall in random, popular, and adversarial sampling settings. Specifically, in the sampling decoding setting, SID surpasses the normal sampling decoding by a large margin in a train-free manner. SID also clearly outperforms CD methods (Dola, ICD, and VCD) because the self-introspective decoding strategy amplifies vision-and-text association hallucinations then subtracts them, rather than coarsely disturbing raw inputs. Additionally, owing to the context and text-aware token selection strategy, SID is more computation-efficient than CD methods, as analyzed in Table 6. Note that beam-search based OPERA (Huang et al., 2024) shows almost no gain in the POPE metric, primarily because answering the binary classification only requires a few tokens and selecting the best beam score in a decoded sequence (=5) brings little improvement.
@@
GPT-4 Assisted Benchmark. While CHAIR and POPE evaluate object-existence-level hallucinations, these metrics are unable to identify other types of hallucination, such as positional, relational, and attribute hallucinations. Therefore, the GPT-4 assisted benchmark (Zhao et al., 2023) utilizes the fine-grained object-level descriptions in the Visual Genome (VG) dataset (Krishna et al., 2017) as ground-truth and relies on the advanced GPT-4 to judge the fine-grained hallucinations and calculate Sentence-level Hallucination Ratio (SHR). Please refer to the Appendix A.3 for detailed implementations. Moreover, we employ n-gram fluency (n = 1 and 2) metrics to measure the smoothness of generated text, and the number of generated words/sentences per image (WPI/SPI) to compare the detailedness of generated texts. As shown in Fig. 7, SID achieves the best results in the SHR metric among the four LVLMs, outperforming others by a clear margin. Regarding the quality of the generated texts, Sampling decoding outperforms ours slightly in terms of 1-gram fluency and WPI. However, compared to other baselines, our approach alleviates hallucinations with minimal sacrifice in text generation quality regarding smoothness and detailness. For instance, OPERA generates text with fewer words and sentences due to penalization of the over-trust mechanism, and VCD impairs text fluency, possibly arising from the holistic and fixed disturbance of contrastive inputs.
MME and MMBench Evaluations. Besides, we test different decoding strategies on two popular LVLMs benchmark: MME (Fu et al., 2023) and MMBench (Liu et al., 2023b). MME comprises ten subtasks to evaluate models’ perceptual capabilities and four subtasks for assessing cognitive abilities, in the form of the yes/no question. MMBench systematically evaluates twenty ability dimensions of LVLMs. We present the results of LLaVA-1.5 7B as a representative in Table 5, SID can maintain and improves the multimodal ability on LVLMs benchmarks. In contrast, other CD methods tend to compromise the general multimodal ability.

### v3

@@ Abstract
Hallucination remains a significant challenge in Large Vision-Language Models (LVLMs). To alleviate this issue, some methods, known as contrastive decoding, induce hallucinations by manually disturbing the raw vision or instruction inputs and then mitigate them by contrasting the outputs of the original and disturbed LVLMs. However, these holistic input disturbances sometimes induce potential noise and also double the inference cost. To tackle these issues, we propose a simple yet effective method named Self-Introspective Decoding (SID). Our empirical investigations reveal that pre-trained LVLMs can introspectively assess the importance of vision tokens based on preceding vision and text (both instruction and generated) tokens. Leveraging this insight, we develop the Context and Text-aware Token Selection (CT2S) strategy, which preserves only the least important vision tokens after the early decoder layers, thereby adaptively amplify vision-and-text association hallucinations during auto-regressive decoding. This strategy ensures that multimodal knowledge absorbed in the early decoder layers induces multimodal contextual rather than aimless hallucinations, and significantly reduces computation burdens. Subsequently, the original token logits subtract the amplified fine-grained hallucinations, effectively alleviating hallucinations without compromising the LVLMs’ general ability. Extensive experiments illustrate that SID generates less-hallucination and higher-quality texts across various metrics, without much additional computation cost111Codes are available at https://github.com/huofushuo/SID.
@@ 1 Introduction
Despite their extraordinary versatility, LVLMs face a significant challenge known as the ‘hallucination’. Concretely, hallucinated texts are fluent and semantically coherent but contain incorrect statements about the given image, e.g., generating irrelevant or meaningless responses, identifying inaccurate colors, numbers, and locations of objects not present in the image (Huang et al., 2024). This flaw in LVLMs poses a significant risk for real-world applications to become trustworthy AI assistants. For instance, in model-assisted computer-aided diagnosis scenarios (Wang et al., 2023), such misinterpretation of medical images could lead to serious medical accidents.
@@ 2 Related Work
Hallucination in Foundation Models. Hallucination, defined as the generation of irrelevant, factually incorrect, or meaningless text in a given context (Rohrbach et al., 2018; Zhang et al., 2024; Guan et al., 2024; Wu et al., 2024c), is a significant bottleneck in current foundation models. This issue can stem from overfitting specific patterns in the training data, a lack of understanding world knowledge, or an inability to effectively contextualize a given input (Ji et al., 2023). In the context of LLMs, hallucinations often manifest as generated content that conflicts with world knowledge or common sense. For LVLMs, the primary concern is whether the generated answer conflicts with the provided images. To mitigate the hallucination issue, several solutions have been proposed, including robust instruction tuning with curated datasets (Lee et al., 2022; Gunjal et al., 2024; Liu et al., 2024a; Zhao et al., 2024; Jiang et al., 2024; Yu et al., 2024b; a; Ma et al., 2024; Yue et al., 2024b), post-hoc utilizing auxiliary analysis networks (Manakul et al., 2023; Zhou et al., 2024; Yin et al., 2023; Chen et al., 2024b; Wu et al., 2024b; Feng et al., 2024), and various decoding strategies (Li et al., 2022; Chuang et al., 2024; Liu et al., 2024c; Leng et al., 2024; Favero et al., 2024; Wang et al., 2024a; Kim et al., 2024; Zhu et al., 2024). However, robust instruction tuning requires massive high-quality datasets and advanced GPU clusters, making it resource-intensive; Post-hoc utilizing auxiliary networks heavily rely on the auxiliary network, leading to high inference costs. As for decoding strategies, representative LVLMs hallucination alleviation methods (Leng et al., 2024; Favero et al., 2024; Wang et al., 2024a) manually disturb raw inputs to induce hallucinations then contrast them to alleviate the issue. However, holistic disturbing raw inputs might bring additional noise during contrastive decoding, and double the inference cost. In this paper, we propose an efficient Self-Introspective Decoding (SID) that induces and then mitigates vision-and-text association hallucination by token-level disturbances, greatly reducing the inference cost.
@@ 3.1 Paradigm of LVLMs Generation
Vision and Language Inputs. The inputs to LVLMs consist of both image () and text (). Generally, the raw images are commonly fed to the visual encoder, and then the cross-model projection module maps vision information into LLMs’ input space, denoted as vision tokens ( is the length of vision tokens). Similarily, text is processed by tokenizer and embedding modules, which is denoted as text tokens ( is length of text tokens). Then, the image () and text () tokens are concatenated as the final input of LLMs.
@@ 3.2 Re-thinking Contrastive Decoding in LVLMs
However, we argue that manually disturbing raw inputs might not trigger the desired hallucinations, while holistic disturbances will bring uncertainty noise that compromises the normal decoding. To validate our claim, we analyze the performances of normal decoding, VCD, and ICD using the POPE (Li et al., 2023d) metric, under both sampling and greedy decoding settings. POPE quantitatively converts the hallucination evaluation into a binary classification problem by using the question format to prompt the model: ‘Is there a <object> in the image?’, with expected answers being ‘Yes’ or ‘No’. From Table 1, under the greedy decoding setting, CD methods improve performance in the adversarial setting, which are more challenging as they prioritize co-occurring confusing objects. CD methods achieve this by exacerbating and subtracting hallucinated concepts from the original distribution. However, in random settings, where objects absent from the image are chosen randomly and are easily recognized, CD methods slightly underperform normal greedy decoding, which indicates that the correct token logit is somewhat compromised during contrastive decoding. In the sampling decoding setting, CD methods clearly outperform the normal sampling decoding. However, CD methods rely on the adaptive plausibility constraint (Eq. 3) to filter out low-probability tokens. Without Eq. 3, CD methods are inferior to normal decoding in both random and adversarial settings, validating that vision-and-text agnostic input distributions induces potential uncertainty noise after Eq. 2 (More validations on other benchmarks are in Table 11.). To address these issues, we propose a decoding strategy named Self-Introspective Decoding (SID). SID adaptively amplifies vision-and-text association hallucinations informed by generated tokens to guide LVLMs in exploring factualness. Details are illustrated in the Sec. 4 and Fig. 2.
@@ 4 Methodology
Analyses. Fig. 3 and 4 preliminarily validate the efficacy of qualitatively. In Fig. 3, the preserved least important tokens mainly reflect areas opposite to the query. For instance, when querying ‘cup’ in Fig. 3 (left), LVLMs focus on ‘cup’ in the foreground, thus preserving background tokens with low . Conversely, LVLMs pay attention to background items when querying ‘couch’. When querying existing items in Fig. 3 (right), vision tokens of unrelated regions are mainly preserved. For open-end generative tasks in Fig. 4, auto-regressive decoded tokens are generated based on preceding vision (), instruction (), and generated text () tokens. The preserved vision tokens are adaptively adjusted according to preceding tokens at each decoding step, primarily focusing on spurious related regions. More quantitative analyses are provided in Appendix A.7, where Table 12 illustrates that vision tokens with high greatly maintain original ability, while tokens with low reach nearly accuracy in binary classification. Above evaluations suggest that Eq. 5 effectively assesses the importance of vision tokens.
We further demonstrate the open-end generated hallucinations induced by ours, Vision Disturbance (VD) (Leng et al., 2024), and Instruction Disturbance (ID) (Wang et al., 2024a) in Fig. 5 and 10. The hallucinations we amplified are more vision-and-text association compared to VD, while LVLMs usually refuse to response to ID. Additionally, we demonstrate the quantitative results for discrimination and generation tasks with VD and ID as inputs in Appendix Table 12. Interestingly, VD and ID do not degrades much especially in discrimination tasks. Experiments imply that disturbed target logits still have the highest probability in most cases, and therefore, contrastive decoded target logits are not enhanced much after Eq. 2, while CD methods are susceptible to potential uncertainty noise.
@@
Based on the above investigations, we argue that to induce context- and text-aware hallucinations for contrastive decoding, only a small percentage of vision tokens with low attention scores should be preserved after the early decoder layers. To validate our claims, we conduct the following experiments: 1) In Vision Encoder (VE), we preserve tokens with low attention values between the token and vision tokens in the penultimate layer, calculated as: . 2) In the LLM decoder, we preserve tokens with low importance score (Eq. 5) across varying layers (). Additionally, we adjust the number of preserved vision tokens. As shown in Fig. 6, firstly, pruning vision tokens in VE based on may not always yield positive gains, as the token lacks information about instructions and generated texts, which are crucial for multimodal understanding. Specifically, pruning all vision tokens resembles VIG (Favero et al., 2024), which contrastively amplifies the vision importance over the language prior by ablating vision inputs. Secondly, aggressive pruning of vision tokens (i.e., 0) after is not optimal. As the ideal induced hallucination distributions are target-co-occurring but suppress target logits, the loss of visual information for subsequent decoding results in visual context diminishing, which can lead to aimless hallucinations due to insufficient grounding in visual information. Thirdly, selecting tokens in the late decoder layers degrades contrastive decoding to normal decoding, as preceding layers of LVLMs already decode and understand multimodal information, which is consistent with LLMs’
@@ 5.1 Experimental Settings
Models and Baselines. We utilize four representative LVLMs: InstructBLIP (Dai & et al, 2023), Shikra (Chen et al., 2023a), LLaVA-1.5 (Liu et al., 2024b) at the 7B scale, and LLaVA-NeXT (Li et al., 2024) at the 8B scale. For detailed model descriptions and results at larger scales, please refer to Appendix A.2 and Table 14 in Appendix A.7. Since our method aims to propose training-free LVLM decoding strategies without the aid of auxiliary networks, we compare six decoding methods: Sampling (Top-p=1), Greedy, Dola (Chuang et al., 2024), and LVLM decoding strategies (VCD (Leng et al., 2024), ICD (Wang et al., 2024a), and OPERA (Huang et al., 2024)). For comprehensive comparisons, we apply VCD and ICD in both sampling (Top-p=1) and greedy decoding settings. Additionally, we validate SID with different decoding strategies, as detailed in Appendix A.7.
@@ 5.2 Evaluation Results
In this section, we follow previous methods (Leng et al., 2024; Wang et al., 2024a; Huang et al., 2024) to evaluate the SID on CHAIR (Rohrbach et al., 2018) and POPE (Li et al., 2023d) metrics. Besides manually designed metrics, we also leverage GPT-4 assisted benchmark (Zhao et al., 2024) to evaluate attribute, location, and relation hallucinations. MME (Fu et al., 2023) and MMBench (Liu et al., 2023b) benchmarks are employed to assess the LVLM’s general ability. Moreover, GPT4-V assisted evaluation (Appendix A.4) on both hallucination alleviation and generated text quality, more general benchmarks evaluation (Appendix A.5) on MathVista (Lu et al., 2024), MMVet (Yu et al., 2024c), LLaVA-Bench (Liu et al., 2023a), MMMU (Yue et al., 2024a), and case studies (Appendix A.6) on LLaVA-Bench-in-the-Wild are in Appendix.
@@
CHAIR and POPE Evaluations. CHAIR (Rohrbach et al., 2018) and POPE (Li et al., 2023d) are quantitative metrics to assess objection hallucinations of VLMs. Please refer to the Appendix A.3 for detailed descriptions of CHAIR and POPE. As for CHAIR, Following (Wang et al., 2024a; Huang et al., 2024; Yue et al., 2024b), we randomly select 500 images from the validation set of the MSCOCO (Lin et al., 2014) dataset and query different LVLMs with the prompt: ‘Please describe this image in detail.’. We set the max new tokens to 512 to generate responses for fair comparisons. As shown in Table 3, our method outperforms other baselines in most cases, validating the effectiveness of SID in open-end generation tasks. Compared to CD methods, SID online adaptively prunes attention-important vision tokens informed by instruction and generated text to induce fine-grained hallucinations for contrastive decoding during open-end text generations. For the POPE metric, which comprises three datasets, we average the results in Table 4. Our method performs best overall in random, popular, and adversarial sampling settings. Specifically, in the sampling decoding setting, SID surpasses the normal sampling decoding by a large margin in a train-free manner. SID also clearly outperforms CD methods (Dola, ICD, and VCD) because the self-introspective decoding strategy amplifies vision-and-text association hallucinations then subtracts them, rather than coarsely disturbing raw inputs. Additionally, owing to the context and text-aware token selection strategy, SID is more computation-efficient than CD methods, as analyzed in Table 6. Note that beam-search based OPERA (Huang et al., 2024) shows almost no gain in the POPE metric, primarily because answering the binary classification only requires a few tokens and selecting the best beam score in a decoded sequence (=5) brings little improvement.
@@
GPT-4 Assisted Benchmark. While CHAIR and POPE evaluate object-existence-level hallucinations, these metrics are unable to identify other types of hallucination, such as positional, relational, and attribute hallucinations. Therefore, the GPT-4 assisted benchmark (Zhao et al., 2024) utilizes the fine-grained object-level descriptions in the Visual Genome (VG) dataset (Krishna et al., 2017) as ground-truth and relies on the advanced GPT-4 to judge the fine-grained hallucinations and calculate Sentence-level Hallucination Ratio (SHR). Please refer to the Appendix A.3 for detailed implementations. Moreover, we employ n-gram fluency (n = 1 and 2) metrics to measure the smoothness of generated text, and the number of generated words/sentences per image (WPI/SPI) to compare the detailedness of generated texts. As shown in Fig. 7, SID achieves the best results in the SHR metric among the four LVLMs, outperforming others by a clear margin. Regarding the quality of the generated texts, Sampling decoding outperforms ours slightly in terms of 1-gram fluency and WPI. However, compared to other baselines, our approach alleviates hallucinations with minimal sacrifice in text generation quality regarding smoothness and detailness. For instance, OPERA generates text with fewer words and sentences due to penalization of the over-trust mechanism, and VCD impairs text fluency, possibly arising from the holistic and fixed disturbance of contrastive inputs.
MME and MMBench Evaluations. Besides, we test on two popular LVLMs’ general ability benchmarks: MME and MMBench. MME comprises ten subtasks to evaluate models’ perceptual capabilities and four subtasks for assessing recognitive abilities in the form of the yes/no question. MMBench systematically evaluates twenty ability dimensions of LVLMs. We present the results of LLaVA-1.5 7B as a representative in Table 5, SID can maintain and improve the multimodal ability on LVLMs benchmarks. In contrast, other CD methods tend to compromise the general multimodal ability. More general benchmark evaluations are in Appendix A.5.